Gucchi:

Which of the following is a function equivalent to f(x) = 3(1.7)4x? f(x) = 25.06x f(x) = 3(8.35)x f(x) = 3(8.25x) f(x) = 487.97x

Gucchi:

Am I right with B?

Gucchi:

@axie

Gucchi:

@tetsxpreme

Timmyspu:

I would believe you are correct. Also just a useful tip always double check your work before you answer the question.

Gucchi:

@timmyspu wrote:
I would believe you are correct. Also just a useful tip always double check your work before you answer the question.
yh i always do

axie:

Yup! Your answer is correct! Why would this be B?

Gucchi:

3(1.7)^4x = 3((1.7)^4)^x = 3(8.35)^x

axie:

Yup!

Gucchi:

thanks